OpenStudy (anonymous):

Match the equation with the correct solution. log 7 24 - log 7 (x + 5) = log 7 8 log 10 x + log 10 (x +21) = 2 log 4 x + log 4 9 = log 4 27 log 2 3x + log 2 5 = log 2 30 2log 10 6 - 1/3 log 10 27 = log 10 X Choose... 12 5 -2 4 3 2 please help

OpenStudy (anonymous):

helppppppppp

OpenStudy (anonymous):

It's simple :)

OpenStudy (anonymous):

can u help

OpenStudy (anonymous):

all you have to do is do "Log of the tall divided by the Log of the small" so the LOG of the bigger number divided by the log of the smaller number is what it equals

OpenStudy (anonymous):

when logs are subtracted, they're divided added, multiplied

OpenStudy (anonymous):

can u show me how to do one

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

choose one

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

ok, 2log 10 6 - 1/3 log 10 27 = log 10 X

OpenStudy (anonymous):

do you have msn or skype i can voice chat and speak easier

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh ok cause my keyboard is broken its a lot harder to type.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

wow your profile pic is amazing you're so hot

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:) <3

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

but im not smart in math :(

OpenStudy (anonymous):

ok back

OpenStudy (anonymous):

I got a 98 in grade 11 math :)

OpenStudy (anonymous):

what grade you in?

OpenStudy (anonymous):

well im taking math on line so i can skip 11 n go to 12 but this math is to hard

OpenStudy (anonymous):

It's simple once you get the hang of it.

OpenStudy (anonymous):

I cant really tyep properly with this keybord its hard its kinda broken

OpenStudy (anonymous):

i could help you lots by voice/cam chat

OpenStudy (anonymous):

i dont have one

OpenStudy (anonymous):

oh. msn?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

nothing at all you have?

OpenStudy (anonymous):

dam i do tutoring in math for 20/hour i could give you lots of free tutoring :)

OpenStudy (anonymous):

do u have facebook

OpenStudy (anonymous):

yeah i do

OpenStudy (anonymous):

but can u please help me first on this math

OpenStudy (anonymous):

ok sure

OpenStudy (anonymous):

ill try with this key board

OpenStudy (anonymous):

2log 10 6 - 1/3 log 10 27 = log 10 X

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[2\log 10 6 - 1/3 \log 10 27 = \log 10 X \]

OpenStudy (anonymous):

2log 10 6 - 1/3 log 10 27 = log 10 X

OpenStudy (anonymous):

log10 6 is .77 times 2 is 1.5

OpenStudy (anonymous):

so 1.5 - 1/3 log 10 27 = log 10 X and log 10 27 is 1.43 times 1/3 is .47 so 1.5 - .47 = log10 x

OpenStudy (anonymous):

1.03 = log10 x

OpenStudy (anonymous):

and do 10^ 1.03 = x

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

and finally we get 10.7 = x

OpenStudy (anonymous):

:) there you go <3

OpenStudy (anonymous):

but thats not an answer

OpenStudy (anonymous):

It's rounding . the answer above is 12

OpenStudy (anonymous):

I rounded it differently so with their method its 12. but its all almost the same basically

OpenStudy (tiffanymak1996):

there's a easier way than to actually calculate them, no calculator needed. 27=3^3, so 1/3log10 27 =3/3log3=log3 2log6 - log3=logx 2(log(2*3) -log3 =logx 2(log2 +log3) -log3 =logx 2log2 +2log3 -log3 =logx 2log2+log3 =logx log4 +log3 = logx log12 =logx x=12

OpenStudy (anonymous):

omg i didnt get it ??? :(

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

exactly what i did. see?

OpenStudy (anonymous):

you just have to simplify. You can put log6 in your calculator. see what you get. and more

OpenStudy (tiffanymak1996):

loga+logb =logab

OpenStudy (anonymous):

yeah i said addition of logs is multiplying

OpenStudy (tiffanymak1996):

well if you put it in a calculator, the answer is not an exact value, and that's not the point of the exercise.

OpenStudy (anonymous):

brendas i can explain easier on facebook if you want

OpenStudy (anonymous):

When you reach grade 12, they'll tell you to put it in a calculator.

OpenStudy (anonymous):

You can multiply the logs when you add them, divide when you subtract

OpenStudy (anonymous):

ok please whats ur email or name so i can look u up

OpenStudy (anonymous):

brendas you can give me your full name ill add you

OpenStudy (tiffanymak1996):

fine, i'm in grade 10, but i'm having maths in stanford, put it in the calculator if you want to.

OpenStudy (anonymous):

brenda sanchez

OpenStudy (anonymous):

oh there are a lot of brenda sanchez on facebook lol which one are you

OpenStudy (anonymous):

tiffany can u help me do another one

OpenStudy (anonymous):

give me ur email on facebook

OpenStudy (anonymous):

markfahmy2009@live.com

OpenStudy (anonymous):

whats yours

OpenStudy (anonymous):

i found u

OpenStudy (anonymous):

oh ok i added you